OpenStudy (anonymous):

Math Help. Algebra 2! Please

OpenStudy (anonymous):

post the problem before asking for help, please

OpenStudy (anonymous):

@Hoa If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the second and third equations, then the first and second equations. 5x + 2y + z = -2 3x + 4y + 3z = 2 -4x - 3y - 3z = 1

OpenStudy (anonymous):

since there is Skullpatrol here, let him help

OpenStudy (anonymous):

you guide me how to solve, does that mean you know how to do? why do you ask for help?

OpenStudy (anonymous):

I have a somewhat understanding of the problem and this is my first time seeing a problem like this. So I'm asking for help as far as like a walk through on how to solve this type of problem

OpenStudy (skullpatrol):

What is the question asking for?

OpenStudy (skullpatrol):



OpenStudy (anonymous):

We have to solve the problem usuing the elimination method. But variable z is eliminated in both the seond and third equation

OpenStudy (skullpatrol):

Yes, that is correct.

OpenStudy (anonymous):

z= -5x-2y-2 right?

OpenStudy (skullpatrol):

Right.

OpenStudy (anonymous):

x=2/21 y= 65/21 z=-182/21

OpenStudy (skullpatrol):

@SerikMB OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (anonymous):

but, the answer is evident, i didn't understand the main question

OpenStudy (anonymous):

Yea I'm confused now

OpenStudy (anonymous):

@skullpatrol: please, be patient to a new comer. they need your confirmation about their solution,

OpenStudy (skullpatrol):

@Hoa Please take over.